# Learning Successor Features with Distributed Hebbian Temporal Memory

### v3

@@ Appendix B Biological Interpretation
Appendix C Encoding Observations
@@
Appendix D Implementation details
@@
In all experiments, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable. For SF formation the agent uses $\gamma=0.8$ and $\gamma=0.9$ for Gridworld and AnimalAI experiments respectively. Maximum SF horizon is set to $50$, but it is usually shorter since we employ early stop condition based on feature variables distribution. That is, if the prediction is close to uniform or the probability of a state with associated positive reward meets a threshold, the planning is interrupted. We use probability thresholds equal to $0.05$ and $0.01$ for Gridworld and AnimalAI respectively.
D.1 LSTM
@@
LSTM baseline was implemented with a single LSTMCell from PyTorch library [52]. It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
@@
For both Gridworld and AnimalAI tasks, LSTM with a hidden state size of 100 is trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
@@
We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups. In all experiments the hidden state represents $10$ categorical variables with $10$ states. That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [40]. The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values. Without symexp the neural network has to represent zero probability with high negative logit values and one-probability with high positive logit values, which is hard to reach with low learning rate and may lead to instabilities. Thus, symexp function makes it faster to reach target values in log space.
D.2 CSCG
@@
CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg). CSCG was trained on buffers of $500$ action steps with $1000$ EM steps until convergence. We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper [53]. This smoothed transition matrix was used as an initialization for the next training batch and for inference. The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy.
@@
In so-called oracle mode, the trajectory buffer is continuously growing, and the transition matrix is trained from scratch every $500$ steps. The buffer is discarded at the environmental change and the transition matrix is reinitialised.
D.3 Episodic Control
@@
Here, we provide an algorithm for SF formation using EC dictionary $\mathcal{D}_{a}$ (see Algorithm 3). In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing faction. Therefore, on each time step, all predicted observations are assumed equally probable. But it’s not a limitation in our set of experiments, since we tested our algorithms in deterministic environments.
@@ D.4 DHTM
Appendix E Experimental Setups
E.1 Gridworld
@@
Samples of setups used in our experiments are depicted in Figure 6. Here, the red colour corresponds to obstacles and the light gray colour to the reward. Floor colours are resampled for each trial, but we ensure that there are no more than five floor colours, including the terminal state. In each episode, the agent and the terminal rewarding state have the same initial positions. The episode ends when the terminal state is entered or the limit of action steps is reached.
Figure 6: Schematic view of POMDP Gridworld setups. The left picture shows the initial setup of the environment, with the reward delivered by the light gray square. The right picture represents the environment after the reward is blocked by obstacles (red squares). Agent’s initial position in the lower left corner.
E.2 AnimalAI
AnimalAI is a testbed inspired by experiments with animals [38]. The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more. We tested our agent in a 10x10-meter room surrounded by walls (see Fig. 7). Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward. The agent and food item have the same fixed initial positions for each episode. The episode ends if the food item is collected or the maximum episode time is reached. To make an agent’s actions more discretized, we repeat its choice three times and skip three frames.
@@ E.2 AnimalAI
Appendix F Additional Experiments
In order to ground our experiments to state-of-the-art model based algorithms, as an additional baseline, we test the smallest version of Dreamer V3 with 18 million parameters. We use the implementation described in Voudouris et al. [54], which is based on Hafner et al. [40] and adapted for AnimalAI.
@@ Appendix F Additional Experiments
Appendix G Glossary

### v4

@@ Appendix C Biological Interpretation
Appendix D Encoding Observations
@@
Appendix E Implementation details
@@
In all experiments in Section 4.1, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable. For SF formation the agent uses $\gamma=0.8$ and $\gamma=0.9$ for Gridworld and AnimalAI experiments respectively. Maximum SF horizon is set to $50$, but it is usually shorter since we employ early stop condition based on feature variables distribution. That is, if the prediction is close to uniform or the probability of a state with associated positive reward meets a threshold, the planning is interrupted. We use probability thresholds equal to $0.05$ and $0.01$ for Gridworld and AnimalAI respectively.
In scaling Gridworld experiments (Appendix H.2), we increase maximum planning horizon up to $100$ steps, and we also had to decrease agent’s temperature to $0.02$ for DHTM and to $0.002$ for EC agent. Discount factor is $0.7$ for DHTM and $0.9$ for EC. We found that, in 10x10 maze, DHTM’s strategy becomes more stable if we increase the positive reward threshold up to $0.4$. EC agent converges faster if we increase its softmax temperature, however, it converges to less optimal trajectory on average.
In the noise tolerance AnimalAI experiment (Appendix H.3), for DHTM we use the same set of agent parameters, except we set the goal probability threshold to $0.5$. We also set the goal detection threshold equal to the mean positive reward since we found that noise variable states mostly get reward assignment slightly above zero, in contrast to signal variables, in which only several states get positive reward assignment.
E.1 LSTM and RWKV
@@
LSTM baseline was implemented with a single LSTMCell from PyTorch library (Paszke et al., 2019). It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
The similar way we implemented RWKV baseline: a single RWKV layer supported by single-layer encoder and decoder. Current public RWKV implementation is a fast evolving framework (Bo, 2021), and for the increased performance it is tightly bound to the offline batch training common for the transformer architectures. In our case we needed a so-called sequential mode for online learning similar to LSTM. Thus, we adapted another public implementation mentioned in the official documentation (RWKV in 150 lines of code).
@@
For both Gridworld and AnimalAI tasks, LSTM and RWKV with a hidden state size of 100 are trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
@@
We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested RNNs into groups. In all experiments the hidden state represents $10$ categorical variables with $10$ states. That is, RNN is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in Dreamer V2/V3 (Hafner et al., 2023). The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values. Without symexp the neural network has to represent zero probability with high negative logit values and one-probability with high positive logit values, which is hard to reach with low learning rate and may lead to instabilities. Thus, symexp function makes it faster to reach target values in log space.
E.2 CSCG
@@
CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg). CSCG was trained on buffers of $500$ action steps with $1000$ EM steps until convergence. We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper (Dedieu et al., 2019). This smoothed transition matrix was used as an initialization for the next training batch and for inference. The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy.
@@
In so-called oracle mode, the transition matrix is trained from scratch every $500$ steps. The buffer is discarded at the environmental change and the transition matrix is reinitialised.
E.3 Episodic Control
@@
Here, we provide an algorithm for SF formation using EC dictionary $\mathcal{D}_{a}$ (see Algorithm 3). In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing fation. Therefore, on each time step, all predicted observations are assumed equally probable. But it’s not a limitation in our set of experiments, since we tested our algorithms in deterministic environments.
@@ E.4 DHTM
Appendix F Experimental Setups
F.1 Gridworld
@@
Samples of setups used in our experiments are depicted in Figure 6. Additional setups for scaling experiments are shown in Figure 7. Here, the red colour corresponds to obstacles (negative numbers) and the light gray colour to the reward (number 4). Letters A and G correspond to the agent’s initial position and goal, correspondingly. Floor colours are resampled for each trial, but we ensure that there are no more than five floor colours, including the terminal state. In each episode, the agent and the terminal rewarding state have the same initial positions. The episode ends when the terminal state is entered or the limit of action steps is reached.
Figure 6: Schematic view of POMDP Gridworld setups. The left picture shows the initial setup of the environment, with the reward delivered by the light gray square (number 4). The right picture represents the environment after the reward is blocked by obstacles (red squares or negative numbers). The letter A depicts the agent’s initial position and the letter G—the goal.
Figure 7: Schematic view of POMDP Gridworld setups for scale experiments. From left to right: 5x5, 7x7, 10x10. The letter A depicts the agent’s initial position and the letter G—the goal.
F.2 AnimalAI
AnimalAI is a testbed inspired by experiments with animals (Crosby et al., 2020). The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more. We tested our agent in a 10x10-meter room surrounded by walls (see Fig. 8). Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward. The agent and food item have the same fixed initial positions for each episode. The episode ends if the food item is collected or the maximum episode time is reached. To make an agent’s actions more discretized, we repeat its choice three times and skip three frames.
@@ F.2 AnimalAI
Appendix G Capacity
In this section, we provide basic theoretical analysis of DHTM’s capacity. Since the hidden state for unpredicted feature state transition is generated randomly, the main source of errors is accidental collision of hidden states. Due to special form of the emission factor (see equation 6), the collision is possible only for the same feature sate. Here we analise the capacity of DHTM for factor grpah depicted in Figure 5. Let’s assume that the stream of feature states is sampled from uniformly distributed categorical random variable. Therefore, the probability to have a collision $p_{c}=\frac{1}{n_{\varphi}n_{v}n_{h}}$, where $n_{v}$—number of hidden variables, $n_{\varphi},n_{h}$ are the number of states for feature and hidden variables correspondingly. Therefore, the probability of having no collisions for a $l$ transitions is equal to $(1-p_{c})^{l}$. Let’s set a threshold $c\in(0,1)$, which determines our confidence in that the sequence of transitions is collision free. Solving the equation $1-c=(1-p_{c})^{l}$ with respect to $l$, we get:
$l=\frac{\log(1-c)}{\log(1-\frac{1}{n_{\varphi}n_{v}n_{h}})}.$ (18)
Using the Taylor expansion for $n_{\varphi}n_{v}n_{h}\to+\inf$, we can see that $l$ grows linearly with the total number of states $n_{\varphi}n_{v}n_{h}$. Therefore, DHTM’s capacity scales linearly with the number of neurons. This, however, can be easily improved if we set additional prior on hidden state activation probability, which would take into account hidden state usage, but we leave it for future research.
Appendix H Additional Experiments
H.1 Dreamer
In order to ground our experiments to state-of-the-art model based algorithms, as an additional baseline, we test the smallest version of Dreamer V3 with 18 million parameters. We use the implementation described in Voudouris et al. (2023), which is based on Hafner et al. (2023) and adapted for AnimalAI.
@@ H.1 Dreamer
H.2 Scalability
In this section, we explore DHTM’s scalability properties and compare them to the EC agent. In RL experiments, like presented in section 4, the number of segments largely depends on the RL algorithm used, its exploration strategy, and how fast it converges to a stable policy. We tested DHTM and EC agents in empty Gridworld of different sizes: 5x5, 7x7 and 10x10, where the agent and the goal are placed in opposite corners (see Fig. 7). As shown in Figure 10A, the DHTM-based agent requires fewer episodes to find the goal; however, the policy becomes less stable as the size of the Gridworld grows. As a consequence of faster convergence, DHTM produces fewer segments than the number of EC’s dictionary entries throughout experiments. It also can be seen that for both DHTM and EC agents, the number of memory units grows non-linearly with the state space size (see Fig. 10B).
Figure 10: Comparison of EC and DHTM-based agents in foraging task in POMDP Gridworld of different sizes: 5x5, 7x7 and 10x10. A. Dynamic of steps required for the agent to reach the goal. B. Number of segments grown (or dictionary entries for EC agent) after $300$ episodes with respect to state space size.
In principle, the number of segments reflects the number of transitions stored, which is proportional to the amount of experience provided to the memory. To show that, we tested DHTM and EC agent in 10x10 empty Gridworld using the same memory parameters as for experiments in section 4.1, but under uniform strategy. That is, agents freely explored the environment during $100$ episodes and at the end the number of segments for DHTM and dictionary size for EC was recorded for each episode length. Figure 11 shows that DHTM’s number of segments and EC’s dictionary size grow linearly with the maximum episode length. In other words, since DHTM lacks generalisation abilities it largely serves as a trajectory buffer, in the same sense as a dictionary in EC.
Figure 11: Comparison of DHTM’s number of segments with EC agent’s dictionary size depending on episode length in Gridworld after uniform exploration during $100$ episodes.
DHTM’s computational complexity (time and memory) of the prediction step grows linearly with respect to the number of segments. Therefore, to use it for huge environments, there is need to prune unnecessary segments or remember only selected episodes. This is a common problem among episodic-like memories. The current approach to form SF also should be adjusted to be scalable, since it requires maximum planning horizon to be equal to the episode length. Among the improved approaches could be to employ n-step TD learning, hierarchical planning or SF caching, to recalculate it less frequently.
H.3 Noise Tolerance
In order to show the key difference between DHTM and its simplified EC version, we tested agents in the same AnimalAI setup as described in 4.2, but we add one noise variable to the VAE encoder. It’s easy to see that the naive mapping from distributed VAE encoding to one categorical variable used for EC doesn’t work in this case, since even identical observations will likely be encoded differently due to the noise variable. In contrast, since DHTM is able to handle distributed representations directly (as the name implies), it still can use non-noise variables independently to learn transitions.
In this experiment, we vary DHTM’s factor size (see Fig. 12). Since factors formed completely random, probability to include the noisy variable increases with factor size and this consequently worsen the performance of the agent. If the factor size is set to the total number of variables, DHTM becomes effectively unfactorised, similarly to its EC counterpart.
Figure 12: Number of action steps the agent takes to reach the goal in a 10x10-meter room in the AnimalAI environment, using DHTM and EC models with a distributed Categorical VAE encoder that includes one additional noisy variable (six variables in total). DHTM’s factor size is varied, determining the number of variables sampled for each factor.
Appendix I Glossary